OpenStudy (anonymous):

Rearrange this formula to make V the subject: A= /sqrt[3]{a} 36V^2

OpenStudy (anonymous):

\[A=\sqrt[3]{36V^2}\]?

OpenStudy (anonymous):

type \sqrt[3]{36v^2}

OpenStudy (anonymous):

/sqrt[3]{36v^2}

OpenStudy (anonymous):

open with \[

OpenStudy (anonymous):

end with \]

OpenStudy (anonymous):

\[sqrt[3]{36V^2}\]

OpenStudy (anonymous):

in any case cube both sides to get \[A^3=36V^2\] then \[V^2=\frac{A^3}{36}\] so \[V=\pm\frac{\sqrt{A^3}}{6}\]

OpenStudy (anonymous):

getting closer you forgot the \ before the sqrt

OpenStudy (anonymous):

whats the plus with the underscore?

OpenStudy (anonymous):

@isuckatmaths i can show you in chat not here because it just renders the math

OpenStudy (anonymous):

or you can right click on what i wrote, choose "show source" and you will see it

OpenStudy (anonymous):

V^2=\frac{A^3}{36}

OpenStudy (anonymous):

start with \[